Justin Barber would like to retire in 25 years, and he estimates that he will need $1,000,000 at this time. He plans to deposit $1,000 a month in a retirement account that pays 8% annual interest.

    1. If Justin makes his first $1,000 deposit next month, how much money will he have by the end of 25 years? Will he have the $1,000,000 that he needs?
    1. How much will Justin have to deposit every month if he wants to have exactly $1 million 25 years from now?
Answer #1
a. Future value =-fv(rate,nper,pmt,pv)
= $ 9,51,026.39
Where,
rate = 8%/12 = 0.006666667
nper = 25*12 = 300
pmt = $                1,000
pv = 0
So,
at the end of 25 years, Justin Barber would have $    9,51,026.39
b. Monthly deposit =-pmt(rate,nper,pv,fv)
= $ 2,628.74
Where,
rate = 8%/12 = 0.006666667
nper = 25*12 = 300
pv = 0
fv = $ 25,00,000.00